How do I find the starndard form of an ellipse with the equation x^2+4y^2-6x-27=0?
Answer by stanbon(48580)

find the standard form of an ellipse with the equation x^2+4y^2-6x-27=0?
------------------------------------------
x^2+4y^2-6x-27=0
x^2-6x+9 + 4y^2 = 27+9
(x+3)^2 + 4y^2 = 36
(x+3)^2/36 + y^2/16 = 1
======================
Cheers,
Stan H.